A body at rest explodes into 2 pieces of masses m1and m2. Show that the ratio of their kinetic energies t1÷t2=m2÷m1

Grade:9

1 Answers

Pranav Jayant Chiddarwar
6 years ago
As no external force is acting on the body wecan conserve momentum. m1v1=m2v2=P
 
t1=1/2 m1v12 = P2/2m1
t2 =1/2m2v22 = P2/2m2
therefore, t1/t2 = m2/m1